An engineer collected average cost and revenue data for Arenson’s FC1 handheld financial calculator. Fixed cost $ 300,000 per year Cost per unit $40 Revenue per unit $70 a. (2 questions) What is the range of the breakeven quantity to variation in the fixed cost from $200,000 to $400,000 per year? Use $50,000 increments. What is the incremental change in the breakeven quantity for each $50,000 change in fixed cost?


1
Expert's answer
2021-10-27T08:30:42-0400

Solution:

Break-even quantity (units) = "\frac{Fixed\;costs}{(Revenue\;per\;unit - Variable\;costs\;per\;unit}"


Break-even quantity when Fixed costs is 200,000 = "\frac{200,000}{(70 - 40)} = \frac{200,000}{30} =" 6,667 units


Break-even quantity when Fixed costs is 250,000 = "\frac{250,000}{(70 - 40)} = \frac{250,000}{30} =" 8,333 units


Break-even quantity when Fixed costs is 300,000 = "\frac{300,000}{(70 - 40)} = \frac{300,000}{30} =" 10,000 units


Break-even quantity when Fixed costs is 350,000 = "\frac{350,000}{(70 - 40)} = \frac{350,000}{30} =" 11,667 units


Break-even quantity when Fixed costs is 400,000 = "\frac{400,000}{(70 - 40)} = \frac{400,000}{30} =" 13,333 units

 

Incremental change (200,000 – 250,000) = 8,333 – 6,667 = 1,666 units

Incremental change (250,000 – 300,000) = 10,000 – 8,333 = 1,667 units

Incremental change (300,000 – 350,000) = 11,667 – 10,000 = 1,667 units

Incremental change (350,000 – 400,000) = 13,333 – 11,667 = 1,666 units